C= 100 + 0.8dy I = 140 G= 200 T= 100+ .2y


1
Expert's answer
2021-03-29T12:04:39-0400

Solution:

To calculate national income (Y):

Y = C + G + 1 + Nx

G = 200, I = 140, T = 100 + 0.2Y

C = 100 + 0.8dy = 100 + 0.8(Y – T) = 100 + 0.8(Y – (100+0.2Y)) = 100 + 0.8Y – 80 + 0.16Y

= 100 – 80 + 0.8Y + 0.16Y = 20 + 0.96Y

C = 20 + 0.96Y

Y = 20 + 0.96Y + 200 + 140

Y = 20 + 200 + 140 +0.96Y

Y – 0.96Y = 360

0.04Y = 360


Y = "\frac{360}{0.04}" = 9000

National Income (Y) = 9,000